2) Two copper plates that are 50 cm by 40 m are sandwiched substance with a dielectric constant of 7.5. What is the p(copper) = 1.72 X 10 Arm D(copper) 8.9 g/cm 5. What is the capacitance of this capacitor? copper) 8.9 g/cm M(Copper) 63.546 g/mole y 40 m are sandwiched between 0.1 mm of a 63 Ao m


two copper plates that are 50cm by 40m are sandwiched between 0.1mm of a substance with a dielectric constant 7.5. what is the capacitance of this capacitor?
Answer #1

capacitance = (dielectric constant*ε0*Area) / (Distance between plates) = k*ε0A/D

= 7.5 * 8.845*10-12*(0.5*0.4)/(0.1*10-3) = 0.13281 µF Answer